OpenStudy (anonymous):

plz help i got 3 questions will medal what is the value of p? 2over 3p=-32 what is the value of y? 3(4y+8) = -12 what is the value of z? 4(z-5) = 11z +50=

OpenStudy (texaschic101):

2/3p = -32 --- divide both sides by 2/3 p = -32/(2/3) p = -32 * 3/2 p = - 96/2 p = - 48 3(4y + 8) = -12 --- divide both sides by 3 4y + 8 = - 4 4y = -4-8 4y = - 12 y = -3 4(z-5) = 11z + 50 -- distribute through the parenthesis 4z - 20 = 11z + 50 4z - 11z = 50 + 20 - 7z = 70 z = - 10